I can not seem to get the answers.
A new cruise line has launched 3 new ships, the pacific, the Mediterranean and the Caribbean. The Caribbean has 16 more deluxe rooms then the pacific. The Mediterranean has 30fewer deluxe rooms then 3 times the number of deluxe rooms on the pacific. Find the number of deluxe rooms for each ship if the total number of deluxe rooms arec701.

Answer by stanbon(75887)

A new cruise line has launched 3 new ships, the pacific, the Mediterranean and the Caribbean. The Caribbean has 16 more deluxe rooms then the pacific. The Mediterranean has 30fewer deluxe rooms then 3 times the number of deluxe rooms on the pacific. Find the number of deluxe rooms for each ship if the total number of deluxe rooms are 701.
---------------------
Equations:
p + m + c = 701
c = p + 16
m = 3p-30
------
Substitute for "c" and for "m" and solve for "p"::
p + 3p-30 + p+16 = 701
5p -14 = 701
5p = 715
p = 143 (# on te Pacific)
c = p + 16 = 159 (# on the Caribbean)
m = 3p-30 = 399 (# on the Mediterranean)
-----------
Cheers,
Stan H.